# UniGLM: Training One Unified Language Model for Text-Attributed Graphs

### v1

@@ 3.2 Adaptive Positive Sample Selection
To extract the structural similar node set $\mathcal{S}_{v}$ of $v$ in Eq. (1) (C1), the conventional protocol aims to randomly sample some nodes from $v$’s neighbors, i.e., nodes directly connected in the original TAG. However, this seemingly intuitive approach is suboptimal within our learning scenarios, given that neighborhood distributions frequently exhibit significant variability both within and across graphs. To effectively consolidate critical structure information across diverse TAGs, we posit that an advanced sampling strategy should account for the following essential factors.
@@
  • Local Structure. Leveraging the local neighborhood structure, i.e., directly connected nodes, constitutes a fundamental design principle of GNNs Kipf and Welling (2016a). To achieve a unified alignment of text attributes and graph structures across various TAGs, the local neighbors of nodes are indispensable.
@@
  • High-Order Structure. Beyond the local structure, high-order neighbors are crucial for the efficacy of graph machine learning Liu et al. (2020); Li et al. (2021), particularly for long-tail nodes Liu et al. (2021).
@@
  • Graph Statistics. Unlike the standard learning paradigm, learning from multiple TAGs necessitates consideration of the unique characteristics inherent to each graph. For example, node degree distributions vary across TAGs, leading to diverse interpretations of hub nodes. Additionally, node status are graph-specific and not directly comparable between TAGs.
@@
Motivated by these observations, we propose an innovative positive sample selection scheme that adaptively sample structurally similar neighbors by considering nodes’ local and high-order neighbors, as well as their unique statuses within each graph. Specifically, we define an Adaptive Positive Sample selection function $\text{AdaPS}()$ as follows:
${S}_{v}=\text{{AdaPS}}(C_{v},W_{v})$ (2)
where the positive samples candidates set is denoted as $C_{v}$ and the corresponding sampling weights is denoted as $W_{v}$. In each iteration, we select positive samples from a candidate set $C_{v}$, which is adaptively chosen to emphasize nodes with more informal, structure-related attributes. The selection is weighted based on the unique statistical characteristics of each graph. Next, we will describe the process of obtaining these candidates and weights.
Given graph ${G}_{i}$ and a central node $v$, positive sample candidates set $\mathcal{C}_{v}$ of $v$ is denoted as: $\mathcal{C}_{v}=\text{Cands}(v,{G}_{i},t)$. We define function Cands() as follows:
$\text{Cands}(v,{G}_{i},t)=\begin{cases}N_{1}(v)+N_{h}(v)&\text{if }\deg(v),\deg(\overline{{G}_{i}})<t\\ N_{1}(v)&\text{otherwise}\end{cases}$ (3)
@@
Here, $\deg(v)$ represents the degree of node $v$ and $\deg(\overline{{G}_{i})}$ is the averaged degree of $G_{i}$. $N_{1}(v)$ and $N_{h}(v)$ denote the first-hop and high-order neighbor set of $v$, respectively. For each central node, we select candidates adaptively by considering both individual node statistics and overall graph metrics. This approach ensures that nodes with local or high-order structure information are chosen as candidates properly for each central node with minimal noise.
To further take personalized node status into consideration, each node within the candidate set for a given central node $v$ is assigned a sampling weight. These weights are calculated using the softmax function applied to the PageRank scores of the candidates as follows.
$W_{v}=\left\{\frac{\exp\left(\frac{\text{PR}(u)}{\tau}\right)}{\sum_{k\in C_{v}}\exp\left(\frac{\text{PR}(k)}{\tau}\right)}\mid u\in C_{v}\right\}$ (4)
where $\text{PR}(u)$ is the PageRank score of node $u$, and $\tau$ is the temperature parameter that controls the concentration of the probability distribution. This strategy ensures that the selection probability of each candidate is proportionate to their personalized node status, thus effectively leveraging their structural prominence within the graph.
@@
Remark. Our design strategically incorporates both local and higher-order structural information through the candidate selection function $C$, enhancing the depth and accuracy of structural insights in TAGs. By further integrating the PageRank scores in our adaptive sampling function $S$, the model prioritizes key nodes based on their centrality, adapting effectively to each graph’s unique statistical characteristics. This comprehensive approach ensures robust performance and superior adaptability across various graph topologies.
3.3 The Lazy Contrastive Module
Another challenge in fine-tuning LMs from multiple TAGs using Eq. (1) is the efficiency problem (C2). Given the constraints of GPU memory, there is a trade-off between the training batch size and the maximum number of positive samples considered. Increasing the batch size can accelerate the training speed by reducing the number of iterations per epoch, which is important given the large scale training nodes in our learning scenarios, yet at the expense of reducing the number of positive samples per node, and vice versa. However, as verified in previous studies Chien et al. (2022); Fang et al. (2024), preserving a sufficient number of structurally similar nodes is crucial to the success of graph contrastive learning.
To address the dilemma, inspired by momentum contrastive He et al. (2020), we introduce a lazy contrastive module by treating positive sample encoding as dictionary look-up operation. Specifically, we establish a dynamic dictionary across various TAGs, which preserves and updates the representations of positive samples on-the-fly using nodes in the batch size, thereby avoiding the need to encode the text attributes of positive samples using LMs during the training. Formally, we rewrite the standard contrastive loss in Eq. (1) to an efficient version, expressed as:
@@ 3.3 The Lazy Contrastive Module
$\begin{split}\mathcal{L}_{v}&=-\sum_{u\in\mathcal{S}_{v}}\log\frac{\exp\left({\text{sim}(f(\mathcal{T}_{v}),\mathbf{y}_{u}})/{\tau}\right)}{\sum\limits_{k\in\mathcal{B}}\exp\left({\text{sim}(f(\mathcal{T}_{v}),f(\mathcal{T}_{k})})/{\tau}\right)}\\ &s.t.\ \ \ \ \mathbf{y}_{u}=\text{LookUp}(\mathbf{E},\text{Idx}(u)).\end{split}$ (5)
@@
Here, $\text{LookUp}(\cdot,\cdot)$ denotes a simple embedding look-up operation based on the embedding table $\mathbf{E}^{n\times d}$ and the index of node $u$, where $n=\sum_{i}^{m}n_{i}$, $\text{Inx}(u)$ depends on both graph index and node index within the graph, and $d$ represents the hidden dimension. Compared to Eq. (1), learning through Eq. (5) is efficient, as the representations of positive samples $\mathbf{y}_{*}$ are obtained via embedding retrieval without the need for explicit text encoding. It is also worth noting that Eq. (5) is memory-efficient since $\mathbf{y}_{*}$ is gradient-free, reducing the abundance of intermediate tensors for gradient calculation. Consequently, a larger batch size can be used to further enhance the training speed. We empirically demonstrate the efficacy of these designs in Table 5 (Appendix). Next, we will show how to effectively implement the lookup operation.
@@
$\begin{split}\mathbf{E}(\text{Idx}(v_{i,j}))&=f(\mathcal{T}_{v_{i,j}}),\quad v_{(i,j)}\in\mathcal{B}.\end{split}$ (6)
Here, only nodes in $\mathcal{B}$ are used to update the embedding table $\mathbf{E}$, which gives rise to the name "lazy", since it utilizes the encoded representations in previous iteration for dictionary updating. In parallel, given the index of positive sample $u_{i,j}$ of node $v_{i,j}$, we extract its hidden representation via simple indexing, i.e., $\mathbf{y}_{u_{i,j}}\mathbf{E}(Idx(u))$, which is LM encoding-free and can accelerate the training speed.
Remark. In contrast to MoCo He et al. (2020), we do not employ an additional momentum LM encoder to update the embedding table $\mathbf{E}$ over time. Instead, we directly utilize the encoded central nodes from previous training steps as the latest representations for updating $\mathbf{E}$. This design not only enhances our training speed, as demonstrated in Appendix Table 5, but also results in notable performance improvements, as empirically verified in Table 1, as it encourages the LM encoder to learn from previous experiences.
@@
Computers SE 57.34±0.47 (+30.47%) 70.22±0.40 (+15.55%) 71.02±0.27 (+15.28%) 70.54±0.23 (+17.11%)
BERT 54.04±0.20 (+38.43%) 66.88±0.24 (+21.32%) 67.25±0.12 (+21.74%) 65.69±0.05 (+25.76%)
GIANT 73.05±0.31 (+2.41%) 80.09±0.16 (+1.31%) 81.03±0.12 (+1.04%) 80.99±0.18 (+2.00%)
PATTON 71.60±0.30 (+4.48%) 78.64±0.14 (+3.18%) 79.98±0.15 (+2.36%) 78.97±0.23 (+4.61%)
MixGIA 65.34±0.16 (+14.49%) 75.13±0.08 (+8.00%) 75.83±0.20 (+7.97%) 75.67±0.25 (+9.17%)
UniGLM 74.81±0.14 81.14±0.19 81.87±0.10 82.61±0.13
Fitness SE 78.07±0.18 (+15.78%) 83.20±0.19 (+9.18%) 83.65±0.17 (+8.98%) 84.26±0.16 (+8.12%)
BERT 74.92±0.26 (+20.65%) 80.77±0.23 (+12.47%) 81.29±0.19 (+12.14%) 80.96±0.49 (+12.52%)
GIANT 89.03±0.07 (+1.53%) 89.63±0.14 (+1.35%) 90.15±0.06 (+1.12%) 90.28±0.10 (+0.91%)
PATTON 89.60±0.22 (+0.88%) 90.03±0.21 (+0.90%) 90.61±0.12 (+0.61%) 90.58±0.22 (+0.57%)
MixGIA 82.83±0.12 (+9.13%) 86.05±0.11 (+5.57%) 86.59±0.16 (+5.28%) 86.63±0.15 (+5.16%)
UniGLM 90.39±0.08 90.84±0.08 91.16±0.11 91.10±0.16
PubMed SE 68.06±2.03 (+20.45%) 73.41±3.02 (+9.06%) 74.25±2.48 (+10.07%) 72.56±1.16 (+11.78%)
BERT 59.79±2.71 (+37.11%) 69.96±2.36 (+14.44%) 63.12±2.43 (+29.48%) 64.34±3.10 (+26.06%)
GIANT 73.18±0.97 (+12.03%) 76.93±0.73 (+4.07%) 74.82±0.65 (+9.24%) 75.54±1.09 (+7.37%)
PATTON 79.64±1.30 (+2.94%) 82.57±0.71 (-3.04%) 80.26±0.95 (+1.83%) 80.58±1.94 (+0.66%)
MixGIA 71.11±2.51 (+15.29%) 76.86±0.61 (+4.16%) 73.71±1.32 (+10.88%) 73.85±1.01 (+9.83%)
UniGLM 81.98±1.32 80.06±1.83 81.73±1.06 81.11±0.69
Photo SE 61.24±0.41 (+25.44%) 71.70±0.16 (+10.88%) 72.14±0.33 (+11.88%) 71.63±0.23 (+12.68%)
BERT 60.03±0.14 (+27.97%) 69.63±0.42 (+14.17%) 70.25±0.36 (+14.89%) 68.79±0.09 (+17.33%)
GIANT 77.43±0.27 (-0.79%) 79.79±0.14 (-0.36%) 81.17±0.23 (-0.57%) 80.69±0.43 (+0.02%)
PATTON 74.97±0.43 (+2.47%) 78.40±0.18 (+1.40%) 78.79±0.01 (+2.44%) 79.79±0.14 (+1.15%)
MixGIA 70.72±0.11 (+8.63%) 75.88±0.10 (+4.77%) 77.34±0.09 (+4.36%) 76.59±0.17 (+5.38%)
UniGLM 76.82±0.29 79.50±0.06 80.71±0.29 80.71±0.02
@@
Table 1: Semi-supervised accuracy results on MLP and state-of-art GNNs with various embeddings for Children, Computers, History, PubMed, and Ogbn-Arxiv datasets. More results can be seen in Appendix.
4 Experiments

### v2

@@ 4.2. Adaptive and Learnable Positive Scheme
To extract the set of similar structural nodes $\mathcal{S}_{v}$ of $v$ in Eq. (1) (C1), the conventional protocol aims to randomly sample some nodes from the neighbors of $v$’, that is, nodes directly connected in the original TAG. However, this seemingly intuitive approach is suboptimal within our learning scenarios, given that neighborhood distributions frequently exhibit significant variability both within and across graphs. To effectively consolidate critical structure information across diverse TAGs, we posit that an advanced sampling strategy should account for the following essential factors.
@@
  • Local Structure. Leveraging the local neighborhood structure, i.e., directly connected nodes, constitutes a fundamental design principle of GNNs (Kipf and Welling, 2016a). To achieve a unified alignment of text attributes and graph structures across various TAGs, the local neighbors of nodes are indispensable.
@@
  • High-Order Structure. Beyond the local structure, high-order neighbors are crucial for the efficacy of graph machine learning (Liu et al., 2020; Li et al., 2021), particularly for long-tail nodes (Liu et al., 2021).
@@
  • Graph Statistics. Unlike the standard learning paradigm, learning from multiple TAGs necessitates consideration of the unique characteristics inherent to each graph. For example, node degree distributions vary across TAGs, leading to diverse interpretations of the hub nodes. Additionally, node status are graph-specific and not directly comparable between TAGs.
4.2.1. Adaptive Positive Sampling
@@
Motivated by these observations, we propose an innovative positive sample selection scheme that adaptively sample structurally similar neighbors by considering nodes’ local and high-order neighbors, as well as their unique statuses within each graph. The selection is weighted on the basis of the unique statistical characteristics of each graph. Given graph ${G}_{i}$ and a central node $v$, positive samples set $\mathcal{S}_{v}$ of $v$ is denoted as: $\mathcal{S}_{v}=\text{AdaPS}(v,{G}_{i},t)$. We define function AdaPS() as follows:
(2) $\small\text{AdaPS}(v,{G}_{i},t)=\begin{cases}N_{1}(v)+N_{h}(v)&\text{if }\deg(v),\deg(\overline{{G}_{i}})<t\\ N_{1}(v)&\text{otherwise}\end{cases}.$
@@
Here, $t$ is the number of positive samples for each central node. $\deg(v)$ represents the degree of node $v$ and $\deg(\overline{{G}_{i})}$ is the average degree of $G_{i}$. $N_{1}(v)$ and $N_{h}(v)$ denote the first-hop and high-order neighbor set of $v$, respectively. If number of nodes in $N_{1}(v)$ exceeds $t$, select the top $t$ nodes with highest Personalized Pagerank Score (Page et al., 1999). For each central node, we select positive samples adaptively by considering both individual node statistics and overall graph metrics. This approach ensures that nodes with local or high-order structure information are chosen as candidates properly for each central node with minimal noise.
@@
Remark: Our design strategically incorporates both local and higher-order structural information through the candidate selection function $C$, enhancing the depth and accuracy of structural insights in TAGs. By further integrating the Personalized PageRank scores in our adaptive sampling function $AdaPS$, the model prioritizes key nodes based on their centrality, adapting effectively to each graph’s unique statistical characteristics. This comprehensive approach ensures robust performance and superior adaptability across various graph topologies.
4.2.2. Contrasting with Informative Positive Samples
Given the sampled positive pool $S_{v}$ of node $v$, one can use Eq. (1) to train the unified PLM across multiple TAGs. However, this approach maybe suboptimal since it treats all positive samples equally. As verified in previous studies (You et al., 2021; Yin et al., 2022), the generation of positive samples are rather important for contrastive learning on graphs. While we do not generate positive samples using explicit graph augmentation strategies, as done in previous efforts (You et al., 2020; Zhang et al., 2024), the adaptive positive generation procedure itself could be regarded as an intuitive implementation of data augmentation, identifying informative positive samples. Thus, the quality of positive samples in $S_{v}$ varies from node to node. In our learning scenario, this may be exacerbated by the inherent heterogeneity of graph types due to the diversity of domains. Therefore, a tailored contrastive objective is needed to boost the performance of our learning from multiple TAGs problem.
To achieve this, we devise a learnable positive generation module, which adaptively selects the most informative positive samples from the pool for loss calculation, by leveraging the personalized node status and domain knowledge of anchor node and corresponding positive candidates. Specifically, let $\mathcal{T}^{D}_{v}$ denote the domain-specific context description of node $v$ including its status and domain knowledge (please refer to Appendix More Experiment Details Section for details), we can rewrite Eq. (1) as follows:
@@ 4.2.2. Contrasting with Informative Positive Samples
(3) $\begin{split}\mathcal{L}_{v}=-\log\frac{\exp\left(\text{sim}(f(\mathcal{T}_{v}),\text{Pos}(\mathcal{T}_{v},S_{v},\mathcal{T}^{D}_{v})))/\tau\right)}{\sum\limits_{k\in\mathcal{B}}\exp\left(\text{sim}(f(\mathcal{T}_{v}),f(\mathcal{T}_{k}))/\tau\right)},\end{split}$
where, $\text{Pos}(\cdot,\cdot,\cdot)$ is a positive generator, which generates an informative positive sample for node $v$ by using the textual attributes of positive candidates in $S_{v}$ and its domain knowledge $\mathcal{T}^{D}_{v}$. In particular, $\text{Pos}()$ adaptively selects the most informative positive candidates to constitute the final positive sample as follows.
(4) $\text{Pos}(\mathcal{T}_{v},S_{v},\mathcal{T}^{D}_{v})=\sum_{u\in S_{v}}\frac{\exp(\mathbf{E}_{pos}^{v}(u))}{\sum_{u^{\prime}}\exp(\mathbf{E}_{pos}^{v}(u^{\prime}))}\cdot f(\mathcal{T}_{u}).$
$\mathbf{E}_{pos}^{v}\in\mathbb{R}^{|S_{v}|}$ is the learnable selection table associated with node $v$, where $\mathbf{E}_{pos}^{v}(u))$ represents the importance of candidate $u$ to $v$. Specifically, $\mathbf{E}_{pos}^{v}(u))$ is initialized as: $\mathbf{E}_{pos}^{v}(u))=\text{sim}(f_{0}(\mathcal{T}_{v}+\mathcal{T}^{D}_{v}),\mathcal{T}_{u})$, and $f_{o}$ stands for the original PLM chosen. By doing this, personalized node characteristics and graph scenarios are incorporated into our learning process.
In the optimization process, we add a regularization term on $\mathbf{E}_{pos}^{v}$ to ensure that the learned importance distribution aligns in general well with the corresponding graph topological structure, expressed as:
(5) $\mathcal{L}_{KL}^{v}=\text{KL}(\text{softmax}(\mathbf{E}_{pos}^{v}),\text{softmax}(\text{Page}^{v})).$
Here, $\text{Page}^{v}\in\mathbb{R}^{|S_{v}|}$ represents the importance vector of node $v$ w.r.t. different nodes in $S_{v}$, where the importance value is estimated by the Personalized PageRank algorithm (Page et al., 1999). KL stands for the Kullback–Leibler divergence between two distributions.
Remark: Our learnable positive generation scheme provides an adaptive way for domain-specific contrastive learning objective. It aligns well with our learning over multiple TAGs scenarios since it can jointly take the graph statistics and node characteristics for positive generation, which excels in identifying the most informative positive sample for node contrasting, as verified in Figure 5.
4.3. The Lazy Contrastive Module
Another challenge in fine-tuning LMs from multiple TAGs using Eq. (1) is the efficiency problem (C2). Given the constraints of GPU memory, there is a trade-off between the training batch size and the maximum number of positive samples considered. Increasing the batch size can accelerate the training speed by reducing the number of iterations per epoch, which is important given the large scale training nodes in our learning scenarios, yet at the expense of reducing the number of positive samples per node, and vice versa. However, as verified in previous studies (Chien et al., 2022; Fang et al., 2024), preserving a sufficient number of structurally similar nodes is crucial to the success of graph contrastive learning.
To address the dilemma, inspired by the momentum contrastive (He et al., 2020a), we introduce a lazy contrastive module by treating positive sample encoding as dictionary look-up operation. Specifically, we establish a dynamic dictionary across various TAGs, which preserves and updates the representations of positive samples on-the-fly using nodes in the batch size, thereby avoiding the need to encode the text attributes of positive samples using LMs during the training. Formally, we rewrite the learnable positive generation process in Eq. (4) to an efficient version, expressed as:
(6) $\begin{split}\text{Pos}(\mathcal{T}_{v},S_{v},\mathcal{T}^{D}_{v})&=\sum_{u\in S_{v}}\frac{\exp(\mathbf{E}_{pos}^{v}(u))}{\sum_{u^{\prime}}\exp(\mathbf{E}_{pos}^{v}(u^{\prime}))}\cdot\mathbf{y}_{u},\\ &s.t.\ \ \ \ \mathbf{y}_{u}=\text{LookUp}(\mathbf{E},\text{Idx}(u)).\end{split}$
Here, $\text{LookUp}(\cdot,\cdot)$ denotes a simple embedding look-up operation based on the embedding table $\mathbf{E}^{n\times d}$ and the index of node $u$, where $n=\sum_{i}^{m}n_{i}$, $\text{Inx}(u)$ depends on both graph index and node index within the graph, and $d$ represents the hidden dimension. The overall loss for node $v$ is computed as $\mathcal{L}_{v}=\mathcal{L}_{\text{vc}}+\alpha\mathcal{L}_{KL}^{v}$, where $\alpha$ is a hyperparameter that adjusts the impact of the KL-divergence loss on the total loss.
@@ 4.3. The Lazy Contrastive Module
Compared to Eq. (4), learning through Eq. (6) is efficient, as the representations of positive samples $\mathbf{y}_{*}$ are obtained via embedding retrieval without the need for explicit text encoding. It is also worth noting that Eq. (6) is memory-efficient since $\mathbf{y}_{*}$ is gradient-free, reducing the abundance of intermediate tensors for gradient calculation. Consequently, a larger batch size can be used to further enhance the training speed. We empirically demonstrate the efficacy of these designs in Table 1. Next, we will show how to effectively implement the lookup operation.
@@
(7) $\begin{split}\mathbf{E}(\text{Idx}(v_{i,j}))&=f(\mathcal{T}_{v_{i,j}}),\quad v_{(i,j)}\in\mathcal{B}.\end{split}$
Here, only nodes in $\mathcal{B}$ are used to update the embedding table $\mathbf{E}$, which gives rise to the name ”lazy”, since it utilizes the encoded representations in previous iteration for dictionary updating. In parallel, given the index of positive sample $u_{i,j}$ of node $v_{i,j}$, we extract its hidden representation via simple indexing, i.e., $\mathbf{y}_{u_{i,j}}\mathbf{E}(Idx(u))$, which is LM encoding-free and can accelerate the training speed.
Remark: In contrast to MoCo (He et al., 2020a), we do not employ an additional momentum LM encoder to update the embedding table $\mathbf{E}$ over time. Instead, we directly utilize the encoded central nodes from previous training steps as the latest representations for updating $\mathbf{E}$. This design not only enhances our training speed, as demonstrated in Table 1, but also results in notable performance improvements, as empirically verified in Table 2, as it encourages the LM encoder to learn from previous experiences.
GIANT PATTON UniGLM w/o E UniGLM
Training time 14.24h 5.4d 63h 13.5h
Inference time 25min 5h15min 20min 20min
Table 1. Training and inference efficiency comparison. For GIANT and PATTON, training time is the sum of time training individual on seven datasets. For UniGLM, it is the time used for training one model with all datasets. For inference time, the time it needs to encode ogbn-arxiv is reported.
@@
Computers SE 57.34±0.47 (+33.24%) 70.22±0.40 (+17.03%) 71.02±0.27 (+16.73%) 70.54±0.23(+17.28%)
BERT 54.04±0.20 (+41.38%) 66.88±0.24 (+22.88%) 67.25±0.12 (+23.27%) 65.69±0.05(+25.94%)
GIANT 73.05±0.31 (+4.59%) 80.09±0.16 (+2.61%) 81.03±0.12 (+2.31%) 80.99±0.18(+2.15%)
PATTON 71.60±0.30 (+6.70%) 78.64±0.14 (+4.50%) 79.98±0.15 (+3.65%) 78.97±0.23(+ 4.76%)
MixGIA 68.83±0.21 (+11.00%) 77.17±0.11 (+6.49%) 76.96±0.26 (+7.72%) 77.22±0.42(+7.14%)
UniGLM 76.40±0.11 82.18±0.08 82.90±0.11 82.73 ± 0.19
Ogbn-Arxiv SE 64.30±0.09 (+15.61%) 71.74±0.29 (+2.82%) 71.49±0.27 (+5.50%) 74.02±0.18(+1.76%)
BERT 66.29±0.20 (+12.14%) 72.63±0.31 (+1.56%) 73.33±0.33 (+2.85%) 72.88±0.39(+3.35%)
GIANT 73.08±0.06 (+1.72%) 73.29±0.10 (+0.64%) 74.59±0.28 (+1.11%) 75.96±0.09(–0.84%)
PATTON 73.47±0.11 (+1.18%) 73.59±0.20 (+0.23%) 75.00±0.16 (+0.56%) 74.08±0.12(+1.67%)
MixGIA 69.30±0.18 (+7.27%) 72.94±0.19 (+1.12%) 73.64±0.20 (+2.42%) 72.39±0.51(+4.05%)
UniGLM 74.34±0.09 73.76±0.12 75.42±0.21 75.32±0.22
Ogbn-Products SE 53.85±0.17 (+43.83%) 70.52±0.51 (+10.55%) 69.13±0.26 (+13.57%) 69.64±0.17(+13.64%)
BERT 67.58±0.28 (+14.60%) 74.77±0.87 (+4.27%) 74.09±0.27 (+5.97%) 74.53±0.26(+6.19%)
GIANT 72.46±0.33 (+6.89%) 69.77±0.42 (+11.74%) 68.69±1.19 (+14.30%) 71.89±0.30(+10.08%)
PATTON 76.42±0.23 (+1.35%) 77.22±0.34 (+0.96%) 77.81±0.58 (+0.90%) 78.48±0.10(+0.84%)
MixGIA 70.44±1.54 (+9.95%) 75.73±0.57 (+2.94%) 75.90±0.84 (+3.44%) 74.88±1.19(+5.69%)
UniGLM 77.45±0.10 77.96±0.22 78.51±0.33 79.14±0.22
Overall SE 60.98±0.24 (+24.22%) 69.66±0.42 (+10.44%) 69.98±0.37 (+10.90%) 70.13±0.35(+10.77%)
BERT 63.66±0.31 (+18.99%) 70.27±0.45 (+9.48%) 70.69±0.40 (+9.79%) 70.22±0.37(+10.62%)
GIANT 73.62±0.28 (+2.89%) 74.56±0.33 (+3.18%) 75.61±0.35 (+2.65%) 76.02±0.34(+2.18%)
PATTON 74.12±0.26 (+2.20%) 75.75±0.31 (+1.56%) 76.80±0.30 (+1.05%) 76.56±0.29(+1.46%)
MixGIA 70.80±0.30 (+6.99%) 74.60±0.34 (+3.12%) 74.92±0.32 (+3.59%) 74.42±0.33(+4.38%)
UniGLM 75.75±0.25 76.93±0.28 77.61±0.27 77.68±0.26
@@
Table 2. Semi-supervised accuracy results on MLP and state-of-art GNNs with various embeddings for History, Computers, Fitness and Ogbn-Arxiv datasets. Top 2 embedding types are bold or underlined. If gap is less than 0.5%, we consider it as both Top 1. More results can be seen in Appendix.
5. Experiments